How do you solve this: 7x +2y = 19
7x + 8y = 13

Found 2 solutions by BryceWalker, Alan3354:
Answer by BryceWalker(4)

Subtract : 7x + 2y = 19
-7x + 8y = 13
this is equal to -6y=5 where y will be equal to -5/6
then let y= -5/6
choose any equation
let have 7x + 2(-5/6) = 19
=> 7x + -5/3 = 19
=> 7x = 19 + 5/3
=> 7x = 62/3
x= 62/3 divided by 7
x= 434/3
on the other hand
x for the other equation would be 79/21
Go for the checking . Hope this answer your problem. by the way, it can be solve by Elimination

Answer by Alan3354(69443)

7x + 2y = 19
7x + 8y = 13
================= Subtract
0x - 6y = 6
y = -1
==============
7x + 2y = 19
7x - 2 = 19
7x = 21
x = 3
--> (3,-1)